OpenStudy (anonymous):

Solve: |2t + 5| = 3 {1, 4} {-1, -4} {-1, 4} {1, -4}

OpenStudy (anonymous):

{-1, -4}

OpenStudy (anonymous):

you need to do 2 problems, 1st using the first part of parenthesis 2(-1)+5 = 3 on the 2nd part, it becomes |2(-4)+5| = |-8+5| = |-3| = 3 the absolute value makes anything inside positive after correct order of operations